OpenStudy (nuccioreggie):

positive two minus negative nineteen equals what value negative twenty-one negative seventeen positive seventeen positive twenty-one

OpenStudy (anonymous):

-7

OpenStudy (paki):

+2- (-19) = 2+19= 21